OpenStudy (anonymous):

Solve the following and show work: 2 log3 (x+2) = log3 9+2

OpenStudy (anonymous):

is this \[2\log_3(x+2)=\log_3(9)+2\]?

OpenStudy (anonymous):

\[\log_3(9)=2\] because \[3^2=9\] so you have \[2\log_3(x+2)=4\] \[\log_3(x+2)=2\] \[x+2=3^2=9\] \[x=7\]

OpenStudy (anonymous):

Yes but no parenthesis around the 9

OpenStudy (anonymous):

that is because it was typeset sloppy

OpenStudy (anonymous):

Thanx for answering I'm still so confused.......not understanding.